OpenStudy (anonymous):

*Will fan and medal!* The time to complete a task varies INVERSELY with the number of people working on the task. It takes 50 people 400 hours to complete a task. How long will it take 80 people to complete the task?

OpenStudy (ribhu):

t = k/p 400 = k/50

OpenStudy (ribhu):

where t = time k = proportionality constant p is no of people required

OpenStudy (ribhu):

k = 20000, putting this in eqn we get t = 20000/80 hrs

OpenStudy (ribhu):

250 hrs @queenleviosa this will solve ur doubt as well as question

OpenStudy (anonymous):

Thanks so much for explaining!

OpenStudy (ribhu):

ok